### v1

# Probing vector chirality in the early Universe

Abstract
We explore the potential of detecting parity violation in primordial vector fossils using late-time galaxy spins. Utilizing $N$-body simulations, we use halo spins as a reliable proxy for galaxy spins to investigate how effectively such primordial vectorial parity asymmetry remains in galaxy spins at low redshifts. We develop a novel approach to generate initial conditions with substantial parity asymmetry, while maintaining the initial matter power spectrum unchanged. From the parity broken initial condition and halos evolved from it, we construct the initial spin and halo spin fields, respectively. Focusing on the helicity of these vector fields, we detect substantial asymmetry in the initial spin field as a consequence of parity violation in the primordial vector fossil. In addition, we discover that over $50\%$ of the primordial asymmetry in the initial spin field remains in the late-time halo spin field on a range of scales. Given the tight correlation between halo spins and observable galaxy spins, we expect to detect the current amplitude of vectorial parity asymmetry potentially up to $16\sigma$-level in observation, when utilizing galaxy samples from DESI BGS. Our findings demonstrate that the primordial imprints of vectorial parity violation persist through non-linear gravitational evolution, highlighting the reliability of galaxy spin as a sensitive probe for testing the vectorial parity-invariance in the early Universe.
preprint: APS/123-QED
Recently, unexpected $3-7\sigma$ detections of parity asymmetry in the observed galaxy distribution were reported [1, 2] focusing on the “helicity” (or handedness) of galaxy quartets in 3-dimension. Although particular care must be taken for accurately modeling systematic uncertainties to conclude such signal to be of primordial origin [2, 3, 4, 5, 6] this motivates further exploration for various types of potential parity violation (PV) in the early Universe and their signatures.
Primary focus has been on the theoretical search for tensorial type, where some inflationary models can produce unequal primordial gravity waves (pGWs) with left- and right-helicities [7, 8, 9, 10, 11, 12]. Thus, detecting either the helical asymmetry between two chiral-pGWs or their imprints on CMB polarization [13, 14, 15, 16] and LSS [17, 18, 19, 20, 21] could confirm the primordial tensorial-PV. Such helical asymmetry imprints would be preferably more detectable in 3-dimensional probes [18] since the helicity information is largely lost in 2-dimension, for example, leaving the detection in the polarized CMB [15, 22, 23, 24] more challenging [c.f. see 25, 26, 27, hinting at the non-vanishing EB-correlation of non-primordial origin, e.g., cosmic birefringence].
Another interesting, but largely unexplored, class of parity violation is vectorial type, possibly generating helically asymmetric vector fossils [17] that may leave imprints in initial vector quantities. Recently, a 3-dimensional estimator “initial spin” (rigorously defined in Eq. 9) – a vector representing the spinning of a mass element torqued by the initial tidal field [28, 29, 30] – was proposed [31, 32] and examined [33] to search for vectorial helical asymmetry. This proposal relies on the two key properties of the initial spin field: 1) a vector field that only consists of left- and right-helical modes, and 2) its strong alignments with “galaxy spin” and “halo spin”, or the angular momentum direction of a galaxy [32, 34] and a dark matter halo [35, 36, 37, 31] at late-time. The latter suggests that the characteristics of the initial spin field will remain in the galaxy spin field. Therefore, detecting helical asymmetry in a galaxy spin field would indicate that the initial spin field is helically asymmetric, evidencing vectorial-PV in the early Universe.
On the observational side, there has not yet been compelling evidence that the galaxy spin field is helically asymmetric; e.g., only discovered were its mildly uneven correlations to the left- and right-helical components of the reconstructed initial spin field [33]. However, it is not straightforward to conclude the absence of primordial vectorial-PV from such a weak-to-none helical asymmetry detection. This is because the primordial helical asymmetry could have been non-negligibly erased during galaxy formation and evolution, leaving only an insignificant signal, or near parity symmetry, in the galaxy spin. This necessitates a further investigation into how effectively the primordial helical asymmetry in the initial spin is carried onto the late-time galaxy spin.
In this Letter, we investigate the detectability of parity violation in a primordial vector fossil, namely “helical” (i.e., rotational) displacement vector field, using galaxy spins. The helical displacement vector field is envisioned to be sourced in the very early universe, e.g., during inflation, and remains as fossils [17]. We emphasize that this vector fossil is not the initial vector mode perturbations, which decay over time. Utilizing numerical N-body simulations, we use halo spins as a reliable proxy for galaxy spins, leveraging the tight correlation between those two spins [38, 39, 37, 34]. In the spirit of effective field theory, without prior assumptions on microscopic physics, we develop a novel way of generating the initial condition with a helically asymmetric initial spin field. We show that broken parity symmetry in the helical displacement fossil efficiently produces strong primordial helical asymmetry without altering the initial matter power spectrum. By analyzing the spins of halos evolved from such initial conditions we demonstrate that such primordial helical asymmetry significantly survives in the final halo spin field.
I Simulation and Initial condition
We construct 500 pairs of N-body simulations – ones with parity symmetric initial condition and the others with parity asymmetric initial condition. The simulations adopt the modified version of the code CUBE [40] to evolve $256^{3}$ dark matter particles within a periodic cubic box of size $L=100~{}h^{-1}{\rm Mpc}$. For numerical convenience, the initial redshift is set $z_{\rm init}=30$. We generate the initial density fluctuation of the standard parity symmetric universe using Zel’dovich approximation [41] where the displacement vector field, $\bm{\Psi}_{\rm sy}$, is curl-free (i.e., irrotational) and its Fourier space representation follows
$\bm{\Psi}_{\rm sy}=\bm{\Psi}_{\rm ZA}=-i\bm{k}\frac{\Phi(\bm{k})}{4\pi Ga^{2}\bar{\rho}}\ ,$ (1)
with the gravitational potential $\Phi$, scale factor $a$, gravitational constant $G$, and mean cosmic density $\bar{\rho}$. On the other hand, the parity asymmetric initial condition is constructed using the modified displacement field, $\bm{\Psi}_{\rm asy}$, including an extra helical displacement, $\bm{\Psi}_{\rm hel}$,
$\bm{\Psi}_{\rm asy}=\bm{\Psi}_{\rm ZA}+\bm{\Psi}_{\rm hel}.$ (2)
The helical displacement field can contain two independent components that are purely right ($\mathcal{R}$)- and left ($\mathcal{L}$)-helical,
$\bm{\Psi}_{\rm hel}$ $=\bm{\Psi}_{\mathcal{R}}+\bm{\Psi}_{\mathcal{L}}=\sum_{\mathcal{H}=\mathcal{R},\mathcal{L}}\Phi_{\mathcal{H}}(\bm{k})\hat{\bm{e}}_{\mathcal{H}}(\bm{k}),$ (3)
where the respective helical displacement field, $\bm{\Psi}_{\mathcal{R}/\mathcal{L}}$, can be constructed as the product of a pseudo-scalar field, $\Phi_{\mathcal{R}/\mathcal{L}}$, and an eigenvector of the curl operator in Fourier space,
$\hat{\bm{e}}_{\mathcal{R}/\mathcal{L}}(\bm{k})=\frac{1}{\sqrt{2}k}\frac{1}{\sqrt{k_{y}^{2}+k_{z}^{2}}}\begin{pmatrix}k_{y}^{2}+k_{z}^{2}\\ -k_{x}k_{y}\pm ikk_{z}\\ -k_{x}k_{z}\mp ikk_{y}\end{pmatrix}\ ,$ (4)
with $k=|\bm{k}|=\sqrt{k_{x}^{2}+k_{y}^{2}+k_{z}^{2}}$. The parity symmetry in the primordial vector fossil is broken if the modified displacement field includes two helical components unequally. To simplify our analysis, we assume $\bm{\Psi}_{\rm hel}$ to be purely right-handed, e.g., $\bm{\Psi}_{\rm asy}=\bm{\Psi}_{\rm ZA}+\bm{\Psi}_{\mathcal{R}}$. Note that both parity symmetric and asymmetric simulations should have identical initial density perturbation in Lagrangian space. This is because the initial density perturbation is solely generated by the same $\bm{\Psi}_{\rm ZA}$ [42], but not by $\bm{\Psi}_{\rm hel}$, as the latter is divergence-free. However, the helical displacement only re-maps the density perturbation at one (Eulerian) position to a different position and plays a pivotal role in breaking parity symmetry in the initial spin vector field, as we will show later.
We now describe the generation of $\Phi$ and $\Phi_{\mathcal{R}}$. We calculate $\Phi$ from the linear matter power spectrum $P(k)$,
$\langle\delta(\bm{k})\delta^{*}(\bm{k}^{\prime})\rangle=(2\pi)^{3}\delta_{D}(\bm{k}-\bm{k}^{\prime})P(k),$ (5)
via the Poisson equation $-k^{2}\Phi=4\pi Ga^{2}\bar{\rho}\delta$, where $\delta$ is the linear density fluctuation. The normalization amplitude is then defined
[…]
and set to $\sigma(z_{\rm init})=0.20$. Note that this corresponds to familiar $\sigma_{8}(z=0)=0.81$, consistent with CMB measurements [43]. Similarly, the pseudo-scalar field $\Phi_{\mathcal{R}}$ is obtained following
$\langle k^{2}\Phi_{\mathcal{R}}(\bm{k})k^{\prime 2}\Phi_{\mathcal{R}}^{*}(\bm{k}^{\prime})\rangle=(2\pi)^{3}\delta_{D}(\bm{k}-\bm{k}^{\prime})P_{\mathcal{R}}(k),$ (7)
where we assume $P_{\mathcal{R}}(k)$ adopts high-$k$ suppression on $P(k)$ as
$P_{\mathcal{R}}(k)=P(k)e^{-k^{2}R_{\rm h}^{2}}.$ (8)
We set $R_{\rm h}=0.6~{}h^{-1}{\rm Mpc}$ in correspondence to the Gaussian radius of a Lagrangian protohalo of $10^{12}~{}h^{-1}{\rm M}_{\odot}$. The normalization is set to $\sigma_{\mathcal{R}}(z_{\rm init})=1.45$ to produce large enough parity asymmetry while maintaining the initial matter power spectrum unchanged. Qualitatively, $\sigma_{\mathcal{R}}$ determines the degree of primordial parity asymmetry, in analogy to non-gaussianity parameter, $f_{\rm NL}$, describing the departure from Gaussianity [44]. A universe without vectorial-PV corresponds to the case with $\sigma_{\mathcal{R}}=0$.
Independent random number sets are used for generating $\Phi$ and $\Phi_{\mathcal{R}}$ to make them uncorrelated. We identify halos using the Friends-of-Friends algorithm and utilize only halos more massive than $10^{12}~{}h^{-1}{\rm M}_{\odot}$. The total number of such halos in the entire 500 simulations of each kind is $N_{\rm halo}\simeq 10^{6}$.
II Initial density field
Figure 1: (Eulerian) Initial matter density fields at $z_{\rm init}=30$ in the parity symmetric (left) and asymmetric (right) simulations sliced in a $0.4~{}h^{-1}{\rm Mpc}$-thickness slab. Density fields are calculated using the could-in-cell (CIC) assignment scheme and smoothed with a Gaussian kernel over $R_{\rm h}=0.6h^{-1}{\rm Mpc}$. Note that they look identical in Lagrangian space.
We begin by visually comparing the initial density fields of the parity symmetric and asymmetric simulations in Fig. 1. It can be seen that the helical displacement field $\Psi_{\mathcal{R}}$ contributes to distorting the standard density field similarly to the gravitationally lensed map of the CMB temperature and polarization [45, 46, 47]. For example, the density structures in the parity asymmetric realization tend to be stretched and/or warped from their counterparts at similar locations in the standard parity symmetric realization. However, unlike CMB maps exhibiting concentric distortion around over/under-densities, the density field deformation due to the helical displacement occurs incoherently.
III Initial and halo spin fields
We now examine the parity symmetry of the initial conditions by measuring the helical asymmetry of the initial spin vector field. The initial spin field, $\bm{J}_{\rm init}$, is calculated on $256^{3}$ grids as a unit vector field constructed from the Hessians of the smoothed initial density, $\rho_{\rm init}^{s}$, and potential field, $\phi_{\rm init}^{s}$, [35, 48, 31, 32, 33],
$\bm{J}_{\rm init}=\left(J_{\rm init,\alpha}\right)\propto\sum_{\beta\gamma\kappa}\epsilon_{\rm\alpha\beta\gamma}\left(\partial_{\beta}\partial_{\kappa}\rho_{\rm init}^{s}\right)\left(\partial_{\gamma}\partial_{\kappa}\Phi_{\rm init}^{s}\right),$ (9)
with the Levi-Civita symbol $\epsilon_{\alpha\beta\gamma}$. The optimal smoothing scale of the Gaussian kernel is set to $R_{\rm s}=2R_{\rm G}$, allowing us to achieve the maximum significance for detecting the primordial vectorial-PV. Note that the initial spin vector field is divergence-free, or purely helical, by construction. On the other hand, the spin vector of a dark matter halo is defined as the unit angular momentum vector of a halo at $z=0$,
$\bm{j}_{\rm halo}\propto\sum_{i}\bm{r}_{p}^{i}\times\bm{v}_{p}^{i},$ (10)
where $\bm{r}_{p}^{i}$ and $\bm{v}_{p}^{i}$ are the relative position and velocity vectors of the $i^{\rm th}$-member particle measured from its halo center. A halo spin field is then constructed by assigning these halo spin vectors to their protohalo centers, $\bm{x}_{\rm ph}$, at $z=z_{\rm init}$,
$\bm{J}_{\rm halo}=\begin{cases}\bm{j}_{\rm halo}&\text{if $\bm{x}=\bm{x}_{\rm ph}^{i}$ for $i^{th}$ halo}\\ \bm{0}&\text{otherwise}.\end{cases}$ (11)
Here, $\bm{x}_{\rm ph}$ is defined as the mass center of halo member particles at their initial positions. We also construct the subsampled initial spin field only using the initial spin vectors at the same protohalo positions $\bm{x}=\bm{x}_{\rm ph}^{i}$, similarly to the halo spin field as in Eq. 11.
IV Measuring Helical asymmetry
We examine the parity symmetry of the initial, subsampled initial, and halo spin fields by measuring the helical asymmetry between their $\mathcal{R}$- and $\mathcal{L}$-handed components. We decompose each of the constructed spin fields into its $\mathcal{R}$- and $\mathcal{L}$-helical fields so that $\bm{J}=\bm{J}_{\mathcal{R}}+\bm{J}_{\mathcal{L}}$. The $\mathcal{R}$$\mathcal{L}$-helical spin field can be computed via
$\bm{J}_{\mathcal{R/L}}=(J_{\mathcal{R/L}})_{\alpha}=\sum_{\beta}\mathbb{P}^{\mathcal{R/L}}_{\alpha\beta}({\bm{k}})J_{\beta}({\bm{k}})$ (12)
using the projection operator defined
$\mathbb{P}^{\mathcal{R/L}}_{\alpha\beta}\equiv\frac{1}{2}\left[(\delta_{\alpha\beta}-\hat{k}_{\alpha}\hat{k}_{\beta})\pm i\sum_{\gamma}\epsilon_{\alpha\beta\gamma}\hat{k}_{\gamma}\right].$ (13)
The helical asymmetry, $\chi$, of a spin field can be then quantified at each wavenumber $k$ following
$\chi(k)\equiv\frac{P_{\mathcal{RR}}(k)-P_{\mathcal{LL}}(k)}{P_{\mathcal{RR}}(k)+P_{\mathcal{LL}}(k)},$ (14)
where $P_{{\mathcal{RR}/\mathcal{LL}}}(k)$ represents the power spectrum of the $\mathcal{R}$$\mathcal{L}$-helical spin field defined as
$\langle\bm{J}_{{\mathcal{R}/\mathcal{L}}}(\bm{k})\cdot\bm{J}^{*}_{{\mathcal{R}/\mathcal{L}}}(\bm{k}^{\prime})\rangle\equiv(2\pi)^{3}\delta_{D}(\bm{k}-\bm{k}^{\prime})P_{{\mathcal{RR}/\mathcal{LL}}}(k)\ .$ (15)
A spin vector field with a perfect helical symmetry satisfies $\chi=0$, whereas $|\chi|$ equals unity when the spin field is maximally asymmetric, or purely $\mathcal{R}$$\mathcal{L}$-helical.
Figure 2: Helical asymmetry spectrum (upper) of initial spin field and the significance of the departure (lower) from the perfect symmetry in the parity symmetric and asymmetric simulations. The vertical line marks a wavenumber corresponding to twice the Gaussian radius of a $10^{12}~{}h^{-1}{\rm M}_{\odot}$ protohalo, $k=1/(2R_{\rm h})$, whereas the horizontal line indicates $\chi/\sigma=50$. Shaded regions represent the standard error of the mean calculated from the 500 realizations, or the total volume of $0.5(h^{-1}{\rm Gpc})^{3}$.
In Fig. 2, we show the helical asymmetry in the full initial spin field in the parity symmetric/asymmetric realizations. We focus only on scales larger than $k=1/(2R_{\rm h})$, below which the helical asymmetry feature in halo spins is likely to be convolved with the halo exclusion effect [49]. In the parity asymmetric realizations, the initial spin field exhibits scale-dependent helical asymmetry that deviates from $\chi=0$, maximally reaching $70\sigma$ detection. Overall, the initial spin field of the parity asymmetric realization is preferentially $\mathcal{R}$-handed, e.g., $\chi>0$. In particular, the helical asymmetry can be as large as $\chi_{\rm peak}\approx 0.18$ indicating that the $\mathcal{R}$-helical mode is stronger by $44\%$ than the $\mathcal{L}$-helical mode. Interestingly, the relative dominance between the two helical modes of the initial spin field is determined to be identical to that of the helical displacement field. On the other hand, as expected, we observe a nearly perfect helical symmetry (i.e., $\chi\approx 0$) on all scales investigated in the parity symmetric simulations.
Figure 3: Same as Fig. 2 but for the subsampled initial spin and late-time halo spin fields. Shaded regions represent the standard errors of the mean using $\sim 10^{6}$ halos. The vertical line again marks $k=1/(2R_{\rm h})$ and grey horizontal bands correspond to $\chi/\sigma\leq 1$ and $2$.
We now compare the late-time halo spin field with the subsampled initial spin field in Fig. 3. Note that we rescaled all helical asymmetry spectra by multiplying the same rescaling factor, which makes the subsampled and full initial spin fields for the parity asymmetric case have identical peak height, e.g., $\chi_{\rm peak}\simeq 0.18$. In the parity symmetric cases, the subsampled initial field shows larger fluctuations around $\chi=0$ than in the full initial spin field. Nonetheless, it is still consistent with the helical symmetry within $1.4\sigma$ significance. Similarly, we do not detect any significant deviation from the perfect symmetry in the halo spin field. The departure at maximum is only about $0.7\sigma$-level across the scales investigated. This indicates that the halo spin field remains parity symmetric if the initial spin field were parity symmetric.
On the other hand, in the parity asymmetric cases, the subsampled initial spin field shows notably large helical asymmetry with scale dependence similar to the full initial spin field. The significance of the asymmetry exceeds $6\sigma$ in the range of scale $0.3\leq k/[h\,{\rm Mpc}^{-1}]\leq 0.7$. For the halo spin field, we again detect considerable helical asymmetry. In particular, the positive departure from the perfect symmetry is detected more than $3\sigma$-significance on scales $0.35\leq k/[h\,{\rm Mpc}^{-1}]\leq 0.75$ and reaches about $5.1\sigma$ at maximum. On these scales, the halo spin field preserves on average more than $50\%$ of the initial helical asymmetry. This clearly demonstrates that the signature of vectorial-PV largely remains even after the gravitational collapse, persisting prominently across these non-linear scales.
V Conclusion and discussion
We assessed the detectability of parity violation in the primordial vector fossil using galaxy spins. Utilizing simulations, we examined how effectively the primordial helical asymmetry is preserved in the late-time galaxy spin field. To this end, we developed a new method to generate initial conditions with parity violating four-point asymmetry as quantified by the initial spin field. Our approach provides a clean channel that effectively embeds parity asymmetry without altering the power spectrum or generating kurtosis via helical remapping, similarly in CMB lensing. This differentiates our approach from the implementation proposed in [50], where larger parity asymmetry inevitably accompanies larger kurtosis and variance.
In the parity asymmetric simulations with $\sigma_{\mathcal{R}}=1.45$, we detected substantial helical asymmetry in the initial spin field, reaching a maximum amplitude of $\chi\simeq 0.18$. More than $50\%$ of such asymmetry on a range of scale survived in the late-time halo spin field and was detected maximally at $5.1\sigma$ significance. This demonstrates that the imprint of primordial vectorial-PV persists prominently in halo spins, only partly reduced by the non-linear gravitational evolution. On the other hand, in the null test (i.e., $\sigma_{\mathcal{R}}=0$), the initial spin was helically symmetric, and so was the halo spin field. Our findings validate the robustness of the halo spin field as tracers of the initial spin field, thereby highlighting its efficacy as a tool for probing the vectorial-PV in the early universe.
Such well-preserved parity asymmetry seen in the halo spins is expected to manifest in observable galaxy spins as well. This is because the spin of baryon components of a galaxy well aligns with its initial spin [34] and host halo spin [38, 51, 39, 34]. Therefore, the initial parity asymmetry should be directly detectable in observation via galaxy spins. If the observed galaxy spin field is helically asymmetric, it would be an unambiguous indication of the vectorial-PV of primordial origin.
Considering the upcoming Y1-data release of the Dark Energy Spectroscopic Instrument (DESI) [52], it presents a timely and compelling opportunity to investigate helical asymmetry in low-redshift galaxy spin fields. The DESI Bright Galaxy Sample [53] will provide high-quality photometric data for over 10 million nearby galaxies, offering precise measurements of position angle, axial ratio, color gradient [54] and winding direction [55], with which 3-dimensional galaxy spins can be unambiguously determined [32]. With such $10^{7}$ galaxy spins, one can potentially reduce the statistical uncertainty in helical asymmetry measurement by a factor of $1/\sqrt{10}$, compared to the present analysis using $10^{6}$ spins. Then, the current parity-violating model with $\sigma_{\mathcal{R}}=1.45$ can be detected up to $16\sigma$ significance in real observation.
Our approach potentially enables us to place observational constraints on theories involving parity-breaking in vector fossils, with the flexibility to accommodate various $\sigma_{\mathcal{R}}$ and $P_{\mathcal{R}}(k)$. An illustrative example closely related to our study would be the generation of the compensated isocurvature perturbations (CIP) [56, 57, 58]. In such scenarios, the helical displacement vector field is introduced to displace dark matter relative to the baryon so that their density perturbations have the same amplitude but opposite signs [59]. If the CIP displacement field exhibits a preferred handedness, i.e., helical asymmetry, the late-time galaxy spin field can be used for constraining these mechanisms.
Finally, exploring how vectorial-PV manifests in various observables presents another intriguing avenue. As proposed in [59], further refinement of the vectorial-PV with the helical displacement field might potentially elucidate the emergence of parity-odd 4-point correlations observed in galaxy distributions [1, 2]. Hence, there is considerable interest in quantifying these 4-point statistics of density fields using our dataset. Another potential signature of vectorial-PV might appear in the intrinsic alignment of galaxy shapes. While the signature of tensorial-PV is expected to appear in the galaxy shape statistics [60, 20, 21], it is yet unexplored for the vectorial-PV. Halo shapes are spin-2 tensors that can be decomposed into tensor and vector modes. Therefore, we might be able to detect asymmetry between the two helical states of these modes. One can also expect to detect a non-zero signal in the inner product between galaxy spin and velocity vectors, as done in [50]. Because this is a parity-odd quantity, its non-zero measurement using our data will confirm the impact of vectorial-PV. Such analyses are left for future exploration.
VI acknowledgments
We thank Donghui Jeong, Eiichiro Komatsu, Oliver Philcox, Shi-Fan Chen, and Tomomi Sunayama for providing valuable feedback on the manuscript. We also thank Academia Sinica Institute of Astronomy and Astrophysics (ASIAA) for hosting the workshop Large-scale Parity Violation Workshop during which this project was advanced. J.S. acknowledges the support by ASIAA. H.R.Y. is supported by National Science Foundation of China grant No. 12173030. T.O. acknowledges support from the Ministry of Science and Technology of Taiwan under grants No. MOST 111-2112-M- 001-061- and No. NSTC 112-2112-M-001-034- and the Career Development Award, Academia Sinica (AS-CDA-108-M02) for the period of 2019-2023. We thank ASIAA for providing computing resources (High Performance Computing Systems).